1. Using the following information, calculate S&G Company's after tax cost of capital.

- The target debt-to-equity is 0.5

- Bonds are currently yielding 10%

- S&G is a constant growth firm that just paid a dividend of $3.00

- S&G common stock sells for $31.50 per share, and has a growth rate of 5%

- Marginal tax rate is 40%

A. 11.0%

B. 12.0%

C. 12.5%
Answer #1
Weighted average cost of capital (WACC) = [(S/S+B)*Rs + (B/S+B)*Rb(1-tc)]
S = equity, B = debt, Rs = Cost of equity, Rb = cost of debt,
tc = corporations tax rate
Rb = .10
tc = .40
According to the dividend growth model.
P0 = D1/(Rs-g)
(price of the stock) P0 = 31.5.
(dividend at the end of year1) D1 = 3*(1.05)
(growth rate of the dividend) g = .05
D1 = 3.15
31.5/(R-.05) = 3.15
R - .05 = 3.15/31.5
Rs = .15
B/S = .5
B = .5S
S/S+B = S/1.5S
S/S+B = .67
S = 2B
B/S+B = B/3B
B/S+B = .33
WACC = [(.67*.15)+((.33)*(.10)*(1-.40))]
WACC = [.1005+((.033)*(.60))]
WACC = [.1005+(.0198)]
WACC = .1203.
S&G company's after tax cost of capital is
B) 12.0%.